# Phase-space localization at the lowest Landau level

### v1

@@ Abstract
We consider bosons with weak contact interactions in a harmonic trap and focus on states at the lowest Landau level. Motivated by the known nontrivial phase-space topography of the energy functional of the corresponding Gross-Pitaevskii equation, we explore Husimi densities of quantum energy eigenstates in the classical phase space of the Schrödinger field. With interactions turned off, the energy levels are highly degenerate and the Husimi densities do not manifest any particular localization properties. With interactions turned on, the degeneracy is lifted, and a selection of energy levels emerges whose Husimi densities are localized around low-dimensional surfaces in the phase space.
Dynamical systems tend to explore all configurations permitted by their conservation laws, in relation to the concept of ergodicity. Localization of the evolution on a surprisingly small set of configurations always stands out as a notable physical signature.
It is especially intriguing when a dynamical system commonly appearing in contemporary theoretical and experimental research manifests such special properties. Here, we turn our attention to nonrelativistic bosons with weak contact interactions confined by an external harmonic potential. This setup is very common in studies of ultracold atomic gases [1, 2]. In a number of situations, including rapid rotation, dynamics of such systems is dominated by the so-called Lowest Landau Level (LLL) sector [2, 3, 4, 5]. We shall focus on this sector, treated at linear order in the small coupling parameter, and observe a subset of quantum energy eigenstates whose corresponding phase-space densities are well-localized.
In discussing phase-space localization, it is important to be clear which phase space we are talking about. Nonrelativistic many-body systems admit equivalent first-quantized and second-quantized representations whose classical phase spaces are very different, resulting in either a mechanical system of particles or a nonrelativistic field theory, respectively. We shall focus on the second-quantized representation and its field-theoretic classical phase space of the corresponding (mean-field) Gross-Pitaevskii equation. It is in this setting that phase-space localization of some of the energy eigenstates in terms of their Husimi densities [6, 7] will be observed. Studies of Husimi densities are a common way to visualize the dynamical content of quantum eigenstates [8, 9, 10, 11, 12, 13, 14]. Our appeal to the phase space of the second-quantized representation is a departure from the common lore, however, and one of our key messages is that this strategy confers many advantages.
Bosons with weak interactions and level splitting.— We start with the Hamiltonian of a self-interacting Schrödinger field $\Psi(x,y)$ in two dimensions in an external isotropic harmonic potential:
@@
The leading corrections $\varepsilon_{I}$ are found by application of the standard first-order degenerate perturbation theory: one evaluates the matrix elements of (3) between the unperturbed eigenstates (5) with the same energy (6) and diagonalizes the resulting matrix. The details of this computation can be found in [15] and we reproduce them in the appendix. Energy levels of (1-3) and related systems have been frequently studied [16, 17, 18, 18, 19, 20, 21, 22, 23, 24, 25], though typically with a focus on small numbers of bosons. We are interested in the sector of the theory spanned by the Fock states (5) where the modes $a_{nm}$ with $n\neq m$ are turned off. This condition defines the LLL sector:
@@
By contrast, in non-LLL states some of the modes with $n\neq m$ have nonzero occupation numbers. Because the LLL modes are located at the edge of the spectrum (there are no modes with $m>n$), it can be shown that $\mathcal{H}_{int}$ has vanishing matrix elements between LLL and non-LLL Fock states. For that reason, one can ignore the non-LLL modes completely for the purpose of computing the first-order energy shifts within the LLL sector. A detailed argument can be found in [15], and is also reproduced in the appendix. At the end of the day, the energy shifts $\varepsilon_{I}$ of LLL states are computed as the eigenvalues of the following LLL Hamiltonian:
@@
This is the quantum version of the classical LLL Hamiltonian describing the weak coupling dynamics of Bose-Einstein condensates in 2d isotropic harmonic traps [26, 27, 28, 29, 30]. The LLL system respects three conservation laws:
@@
$N$ is the total particle number, $M$ is the angular momentum, while $Z$ boosts the center-of-mass angular momentum [31].
We arrived at the Hamiltonian (9) starting from the problem of finding energy corrections (7) to the degenerate levels (6). Such problems amount to diagonalizing finite-dimensional matrices made of the matrix elements of (3) between states within the same degenerate unperturbed level. Even though $\mathcal{H}_{LLL}$ acts on an infinite-dimensional Fock space, its matrix elements split correspondingly into finite-dimensional blocks acting on the Fock states within the same unperturbed degenerate level [32]. Indeed, $\mathcal{H}_{LLL}$ has vanishing matrix elements between Fock states with different values of $N$ and $M$. But for any given values of $N$ and $M$, there is only a finite number of Fock vectors available [33]. One can then straightforwardly diagonalize $\mathcal{H}_{LLL}$ within any such $(N,M)$-block, and obtain the corresponding energy shifts for the degenerate level with $N$ particles, $M$ units of angular momentum, and unperturbed energy (6), also equal to $M$.
The ladder states.— The energy shifts (7), obtained by diagonalizing $\mathcal{H}_{LLL}$ restricted to a particular $(N,M)$-block, are in general complicated real numbers. There is a sprinkling of rational values within this spectrum, however, suggestive of some solvable structure. This structure has been manifested in [34], resulting in a family of explicit, exact eigenstates. It is for these special states that we find strong phase-space localization properties.
We refer the reader to [34] for technical derivations, and simply quote here the key results that can be independently verified by a variety of analytic or numerical means [35]. Within each $(N,M)$-block with $2\leq M\leq N$, the state
@@
where the function $|p(t)|$ turns out to be time-periodic. In [34], it was shown how to form coherent-like combinations of the quantum states (12) that precisely reproduce (16) in the semiclassical limit.
The LLL energy functional.— A key ingredient in the derivation of the states (12) is the following operator [34]:
@@
By writing $B$ in terms of creation-annihilation operators, one notices that it annihilates the state $|N-M,M,0,0,0\dots\rangle$, and consequently also the states (12) and all their $Z$-images (note that $B$ commutes with $Z$ and $Z^{\dagger}$). The condition of being annihilated by $B$ and $Z^{\dagger}$ is an effective way to define the ladder states (12), and it also fixes their energies as (13-14).
@@
which was subsequently proved in [36]. It was moreover shown that the configurations saturating the bound (18) are precisely of the form (16). Since the configurations in (16) are parametrized by 3 complex numbers $a$, $b$ and $p$, the constant energy surface for the energy value saturating the bound (18) has at most 6 real dimensions. This is in striking contrast with generic constant energy surfaces in the infinite-dimensional phase space of $\alpha_{n}$, which are themselves infinite-dimensional. Turning back to the corresponding quantum system, the ladder state energies are slightly below [37] the classical bound (18) and so they are very close to the bottom of the peculiar low-dimensional phase-space valley dictated by (18). One thus expects that the ladder states will be very constrained in the way they explore the infinite-dimensional phase space and will be localized around low-dimensional surfaces. In the remainder of our treatment, we shall show quantitatively that this indeed happens.
Phase-space localization of Husimi densities.— Our strategy will rely on Husimi densities [6, 7], widely used to develop phase-space pictures of quantum eigenstates [8, 9, 10, 11, 12, 13, 14]. For a given state $\psi$, the Husimi density is defined by [38]
@@
These distributions peak at $|\alpha_{i}|^{2}=\eta_{i}$ and are constant under independent phase shifts of $\alpha_{i}$. The Husimi density of a typical Fock state is localized around a high-dimensional torus in phase space [39]. For a random linear combination of Fock states at fixed (large) $N$ and $M$, the distribution typically spreads even further over a high-dimensional region constrained only by the conservation of $N$ and $M$.
When $g>0$, the degeneracies are lifted, and the new nondegenerate eigenstates are determined by the LLL Hamiltonian (9). The Husimi distribution of a Hamiltonian eigenstate is generically expected to decay exponentially fast away from the classical phase-space hypersurface defined by the constant values of the classical conserved quantities corresponding to the chosen eigenstate [41, 40]. For the LLL system, this hypersurface is generally codimension 4 as energy eigenstates are simultaneous eigenstates of the operators $\mathcal{H}_{LLL}$, $N$, $M$ and $ZZ^{\dagger}$. There is one further salient feature of the Husimi densities related to its symmetries. Symmetries of a Hamiltonian do not immediately translate into symmetries of the Husimi distributions. Indeed, the Hamiltonian itself induces a time-translation symmetry, but advancing along classical trajectories does not keep the Husimi density constant. However, specifically for conservation laws of the form $\hat{h}=\sum_{n}h(n)a^{\dagger}_{n}a_{n}$, the associated symmetry transformations $\alpha_{n}\to e^{i\theta h(n)}\alpha_{n}$ leave the Husimi density invariant. We provide a proof of this statement in the appendix. Since $N$ and $M$ in (11) are precisely of this form, the Husimi densities are constant over the two-dimensional tori
@@
Ladder states have energies just below the classical bound (18), and correspondingly we do not expect their Husimi densities to stray far away from the classical configurations (16) that saturate this bound, parameterized by 6 real numbers coming from the complex variables $a$, $b$, $p$. The accessible configurations are further constrained by the 4 independent conservation laws ($N$, $M$, $Z$ and $\bar{Z}$) leaving a two-dimensional surface of accessible configurations [42]. This is the minimal number of dimensions one could in principle have, given the constancy of Husimi densities over the tori (22).
To confirm this picture, we investigate numerically the phase-space localization properties of ladder states, and contrast their Husimi distributions with those of generic states [43]. Our approach consists in exploring the phase space close to the shell $M_{\psi}$ of codimension 4 defined by classical conserved quantities being equal to the corresponding quantum numbers of the state $\psi$ [44],
@@
since the Husimi function should be localized there for all states. We shall start by identifying a configuration $\alpha_{\max}$ where the Husimi density takes a particularly large value. This value will have to be replicated over the entire two-dimensional surface $\Theta(\alpha_{\max})$. We will then see that the Husimi function decays exponentially away from $\Theta(\alpha_{\max})$ for ladder states, while for generic states we encounter a landscape of appreciable Husimi density values spread out over $M_{\psi}$.
We need to find a sample cloud of configurations where the Husimi distribution values are substantial, and compare the resulting geometries of Husimi distributions for ladder and non-ladder states. Due to the high dimensionality of the phase space, judicious choices should be made in designing this sampling. If the cloud is not sufficiently spread out, it will overlook the key features of Husimi density in areas falling outside its coverage. If the cloud is too dispersed, an unrealistically large number of sample points will be required in a high-dimensional space to ever hit appreciable values of the Husimi density. Thus, we must generate a cloud that restricts the search to a useful region of phase space, without being too restrictive. In practice, after constructing, using optimization algorithms, the phase-space point $\alpha_{\max}$ where the Husimi distribution peaks, we use it as the starting value for a Metropolis-Hastings sampling run that generates a cloud of points. To make the procedure effective, as the target probability distribution for this run we choose a Gaussian decaying away from the classical hypersurface (23), reminiscent of the formulas seen in [41, 40],
@@
The subscript ‘$\max$’ refers to the value of the corresponding functional on $\alpha_{\max}$. For concreteness, we focus on a large block of dimension $5604$, with $N=70$ and $M=30$, and restrict our attention to eigenstates with the center-of-mass at rest, annihilated by $Z^{\dagger}$. The relevant ladder state (12) is the lowest energy state in the block. For our choice of a generic state, we pick energy level $5$ above the ground state. To compare the two states, we plot the Husimi function values of the sampled points against their distance $d$ to the two-dimensional surface $\Theta(\alpha_{\max})$ [45]. The results are displayed in Fig. 1 for the two chosen states. Some further numerical examples are given in the appendix, and they are consistent with the general picture presented here.
For the ladder state, Fig. 1 shows an exponential correlation between the Husimi value and the distance away from $\Theta(\alpha_{\max})$. This reflects exponential decay away from the classically permitted region of motion, which coincides with $\Theta(\alpha_{\max})$ in this case. On the other hand, for
@@
Figure 1: Husimi function values sampled over a cloud of phase-space points, plotted against their distance $d$ to the two-dimensional surface $\Theta(\alpha_{\max})$ where the largest Husimi function value lies. Top: for the ladder state with $|Z|=0$; bottom: for a non-ladder state with $|Z|=0$, both in the Hilbert space block with $N=70$ and $M=30$. The Gaussian distribution used in the sampling algorithm has mean values determined by the classical conservation laws at $\alpha_{\max}$, with $(N_{\max},M_{\max},Z_{\max},H_{\max})=(70,30,0,1925)$ (top) and $(N_{\max},M_{\max},Z_{\max},H_{\max})=(70,30,0,1947.04)$ (bottom), obtained by global optimization of the Husimi density starting from a randomly picked $\alpha_{n}$. The standard deviations of the Gaussian distribution used for sampling are set to $5\sqrt{2}\times(\sigma_{N},\sigma_{M},\sigma_{|Z|^{2}},\sigma_{H})=(1,1,N,N)$ for both cases. For the non-ladder state, we further perform local optimization of the Husimi density for randomly selected points in the upper right area of the plot, and display the distance from the resulting local maximum to the initial two-dimensional surface, indicated by the color of the selected points. These results show that the Husimi density of the ladder state is localized on $\Theta(\alpha_{\max})$, while the region of high Husimi density spreads beyond $\Theta(\alpha_{\max})$ for generic states, with many distinct local maxima decorating the landscape.
the generic state, the bottom plot in Fig. 1 shows a protrusion of points in the top-right direction where the Husimi function is large while being far from the two-dimensional surface $\Theta(\alpha_{\max})$.

### v2

@@ Abstract
We consider bosons with weak contact interactions in a harmonic trap and focus on states at the lowest Landau level. Motivated by the known nontrivial phase-space topography of the energy functional of the corresponding Gross-Pitaevskii equation, we explore Husimi distributions of quantum energy eigenstates in the classical phase space of the Schrödinger field. With interactions turned off, the energy levels are highly degenerate and the Husimi distributions do not manifest any particular localization properties. With interactions turned on, the degeneracy is lifted, and a selection of energy levels emerges whose Husimi distributions are localized around low-dimensional surfaces in the phase space.
I Introduction
Dynamical systems tend to explore all configurations permitted by their conservation laws, in relation to the concept of ergodicity [1, 2, 3]. Localization of the evolution on a surprisingly small set of configurations always stands out as a notable physical signature.
It is especially intriguing when a dynamical system commonly appearing in contemporary theoretical and experimental research manifests such special properties. Here, we turn our attention to nonrelativistic bosons with weak contact interactions confined by an external harmonic potential. This setup is very common in studies of ultracold atomic gases [4, 5]. In a number of situations, including rapid rotation, dynamics of such systems is dominated by the so-called Lowest Landau Level (LLL) sector [5, 6, 7, 8]. We shall focus on this sector, treated at linear order in the small coupling parameter, and observe a subset of quantum energy eigenstates whose corresponding phase-space distributions are well-localized.
In discussing phase-space localization, it is important to be clear which phase space we are talking about. Nonrelativistic many-body systems admit equivalent first-quantized and second-quantized representations whose classical phase spaces are very different, resulting in either a mechanical system of particles or a nonrelativistic field theory, respectively. We shall focus on the second-quantized representation and its field-theoretic classical phase space of the corresponding (mean-field) Gross-Pitaevskii equation. It is in this setting that phase-space localization of some of the energy eigenstates in terms of their Husimi distributions [9, 10] will be observed. Studies of Husimi distributions are a common way to visualize the dynamical content of quantum eigenstates [11, 12, 13, 14, 15, 16, 17]. Our appeal to the phase space of the second-quantized representation is a departure from the common lore, however, and one of our key messages is that this strategy confers many advantages.
II Bosons with weak interactions and level splitting
We start with the Hamiltonian of a self-interacting Schrödinger field $\Psi(x,y)$ in two dimensions in an external isotropic harmonic potential:
@@ II Bosons with weak interactions and level splitting
The leading corrections $\varepsilon_{I}$ are found by application of the standard first-order degenerate perturbation theory: one evaluates the matrix elements of (3) between the unperturbed eigenstates (5) with the same energy (6) and diagonalizes the resulting matrix. The details of this computation can be found in [18] and we reproduce them in Appendix A. Energy levels of (1-3) and related systems have been frequently studied [19, 20, 21, 21, 22, 23, 24, 25, 26, 27, 28], though typically with a focus on small numbers of bosons. We are interested in the sector of the theory spanned by the Fock states (5) where the modes $a_{nm}$ with $n\neq m$ are turned off. This condition defines the LLL sector:
@@
By contrast, in non-LLL states some of the modes with $n\neq m$ have nonzero occupation numbers. Because the LLL modes are located at the edge of the spectrum (there are no modes with $m>n$), it can be shown that $\mathcal{H}_{int}$ has vanishing matrix elements between LLL and non-LLL Fock states. For that reason, one can ignore the non-LLL modes completely for the purpose of computing the first-order energy shifts within the LLL sector. A detailed argument can be found in [18], and is also reproduced in Appendix A. At the end of the day, the energy shifts $\varepsilon_{I}$ of LLL states are computed as the eigenvalues of the following LLL Hamiltonian:
@@
This is the quantum version of the classical LLL Hamiltonian describing the weak coupling dynamics of Bose-Einstein condensates in 2d isotropic harmonic traps [29, 30, 31, 32, 33]. The LLL Hamiltonian (9) commutes with the operators
${N}=\sum{a}^{\dagger}_{k}{a}_{k},\,\,{M}=\sum k\,{a}^{\dagger}_{k}{a}_{k},\,\,{Z}=\sum\sqrt{k+1}\,a^{\dagger}_{k+1}a_{k}$ (11)
and $Z^{\dagger}$. The nonvanishing commutators of these operators are
@@
Here, $N$ is the total particle number, $M$ is the angular momentum, while $Z$ boosts the center-of-mass angular momentum. (The center-of-mass motion in harmonic potentials separates from other degrees of freedom, and the center-of-mass performs simple harmonic oscillations. This is an example of a ‘breathing mode.’ General implications of breathing modes for weakly nonlinear dynamics are treated in [34].) Quantum numbers correspond to mutually commuting operators, which can be taken as $\mathcal{H}_{LLL}$, $N$, $M$ and $ZZ^{\dagger}$.
We arrived at the Hamiltonian (9) starting from the problem of finding energy corrections (7) to the degenerate levels (6). Such problems amount to diagonalizing finite-dimensional matrices made of the matrix elements of (3) between states within the same degenerate unperturbed level. Even though $\mathcal{H}_{LLL}$ acts on an infinite-dimensional Fock space, its matrix elements split correspondingly into finite-dimensional blocks acting on the Fock states within the same unperturbed degenerate level. Indeed, $\mathcal{H}_{LLL}$ has vanishing matrix elements between Fock states with different values of $N$ and $M$. For any given values of $N$ and $M$, there is only a finite number of Fock vectors available. (These vectors are indexed by partitions of $M$ into at most $N$ parts.) One can then straightforwardly diagonalize $\mathcal{H}_{LLL}$ within any such $(N,M)$-block, and obtain the corresponding energy shifts for the degenerate level with $N$ particles, $M$ units of angular momentum, and unperturbed energy (6), also equal to $M$. (The property of splitting into independent finite-dimensional block is shared by all Hamiltonians of the form (9), for any values of the mode couplings $C_{nmkl}$. Systematic studies of such Hamiltonians were initiated in [35].)
III The ladder states
The energy shifts (7), obtained by diagonalizing $\mathcal{H}_{LLL}$ restricted to a particular $(N,M)$-block, are in general complicated real numbers. There is a sprinkling of rational values within this spectrum, however, suggestive of some solvable structure. This structure has been manifested in [36], resulting in a family of explicit, exact eigenstates. It is for these special states that we find strong phase-space localization properties.
We refer the reader to [36] for technical derivations, and simply quote here the key results that can be independently verified by a variety of analytic or numerical means. (Considerations of [36] have been generalized to relativistic field systems in [37].) Within each $(N,M)$-block with $2\leq M\leq N$, the state
@@ III The ladder states
where the function $|p(t)|$ turns out to be time-periodic. In [36], it was shown how to form coherent-like combinations of the quantum states corresponding to (15) that precisely reproduce (17) in the semiclassical limit. In this paper, we will be mainly interested in coherent-like combinations of the lowest ladder states (13), which correspond to a subclass of the solutions (17) with $Z=0$ [31, 36].
IV The LLL energy functional
A key ingredient in the derivation of the states (13) is the following operator [36]:
@@ IV The LLL energy functional
By writing $B$ in terms of creation-annihilation operators, one notices that it annihilates the state $|N-M,M,0,0,0\dots\rangle$, and consequently also the states (13) and all their $Z$-images (note that $B$ commutes with $Z$ and $Z^{\dagger}$). The condition of being annihilated by $B$ and $Z^{\dagger}$ is an effective way to define the ladder states (13), and it also fixes their energies as (14). Acting on (13) with $Z$ one obtains further ladder states with energies (15). We shall focus below on the lowest ladder states (13) in each block, since the remaining ones are obtained from those by boosting the center-of-mass motion.
@@
which was subsequently proved in [38]. It was moreover shown that the configurations saturating the bound (19) are precisely of the form (17). Since the configurations in (17) are parametrized by 3 complex numbers $a$, $b$ and $p$, the constant energy surface for the energy value saturating the bound (19) has at most 6 real dimensions. This is in striking contrast with generic constant energy surfaces in the infinite-dimensional phase space of $\alpha_{n}$, which are themselves infinite-dimensional. Turning back to the corresponding quantum system, the ladder state energies are slightly below the classical bound (19) and so they are very close to the bottom of the peculiar low-dimensional phase-space valley dictated by (19). (The difference between the quantum and classical bound $B\geq 0$ is due to ordering ambiguities. The ladder state energies slightly violate the classical bound due to the normal-ordered form of the quantum Hamiltonian, which effectively involves a subtraction of some amount of energy.) One thus expects that the ladder states will be very constrained in the way they explore the infinite-dimensional phase space and that they will be localized around low-dimensional surfaces. In the remainder of our treatment, we shall show quantitatively that this indeed happens.
V Phase-space localization of Husimi distributions
Our strategy will rely on Husimi distributions [9, 10], widely used to develop phase-space pictures of quantum eigenstates [11, 12, 13, 14, 15, 16, 17]. For a given state $\psi$, the Husimi distribution is defined by
@@ V Phase-space localization of Husimi distributions
Husimi distributions are closely related to the more familiar Wigner phase-space distributions. One issue with the latter is that they are not necessarily positive and do not provide genuine probability distributions. An integral convolution of the Wigner function with a Gaussian filter results in the Husimi distribution, which is by construction always positive, see [39].
@@
These distributions peak at $|\alpha_{i}|^{2}=\eta_{i}$ and are constant under independent phase shifts of $\alpha_{i}$. The Husimi distribution of a typical Fock state is localized around a high-dimensional torus in phase space. (These are just the Kolmogorov tori of the trivially integrable linear noninteracting system.) For a random linear combination of Fock states at fixed (large) $N$ and $M$, the distribution typically spreads even further over a high-dimensional region constrained only by the conservation of $N$ and $M$.
When $g>0$, the degeneracies are lifted, and the new nondegenerate eigenstates are determined by the LLL Hamiltonian (9). The Husimi distribution of a Hamiltonian eigenstate is generically expected to decay exponentially fast away from the classical phase-space hypersurface defined by the constant values of the classical conserved quantities corresponding to the chosen eigenstate [41, 40]. For the LLL system, this hypersurface is generally codimension 4 as energy eigenstates are simultaneous eigenstates of the operators $\mathcal{H}_{LLL}$, $N$, $M$ and $ZZ^{\dagger}$. There is one further salient feature of the Husimi distributions related to its symmetries. Symmetries of a Hamiltonian do not immediately translate into symmetries of the Husimi distributions. Indeed, the Hamiltonian itself induces a time-translation symmetry, but advancing along classical trajectories does not keep the Husimi distribution constant. However, specifically for conservation laws of the form $\hat{h}=\sum_{n}h(n)a^{\dagger}_{n}a_{n}$, the associated symmetry transformations $\alpha_{n}\to e^{i\theta h(n)}\alpha_{n}$ leave the Husimi distribution invariant. We provide a proof of this statement in Appendix B. Since $N$ and $M$ in (12) are precisely of this form, the Husimi distribution remains constant over the surface of the two-dimensional torus defined by
@@
Ladder states have energies just below the classical bound (19), and correspondingly we do not expect their Husimi distributions to stray far away from the classical configurations (17) that saturate this bound, parameterized by 6 real numbers coming from the complex variables $a$, $b$, $p$. The accessible configurations associated to a quantum state $|\psi\rangle$ are further constrained by its quantum numbers, which are the eigenvalues of $\mathcal{H}_{LLL},N,M$ and $ZZ^{\dagger}$. For a generic ladder state, only three of these are independent, yielding three constraints. In the case of the lowest ladder state where $ZZ^{\dagger}=0$, however, the corresponding classical constraint is $|Z|^{2}=0$, which is equivalent to $Z=0$. This results in a total of four constraints and leaves a two-dimensional surface of accessible configurations. This is the minimal number of dimensions one could in principle have, given the constancy of Husimi distributions over the tori (23).
To confirm this picture, we investigate numerically the phase-space localization properties of ladder states, and contrast their Husimi distributions with those of generic states [42]. Our approach consists in exploring the phase space close to the shell $M_{\psi}$ of codimension 4 defined by classical conserved quantities being equal to the corresponding quantum numbers of the state $\psi$,
@@
since the Husimi function should be localized there for all states. We shall start by identifying a configuration $\alpha_{\max}$ where the Husimi distribution takes a particularly large value. This value will have to be replicated over the entire two-dimensional surface $\Theta(\alpha_{\max})$. We will then see that the Husimi function decays exponentially away from $\Theta(\alpha_{\max})$ for ladder states, while for generic states we encounter a landscape of appreciable Husimi distribution values spread out over $M_{\psi}$. (Note that the classical energy $H$ defining $M_{\psi}$ differs slightly from the quantum energy of $\psi$ due to ordering ambiguities. The relative discrepancy vanishes in the semiclassical limit $N\rightarrow\infty$. In practice, we determine the classical values defining the shell $M_{\psi}$ by evaluating the conservation laws at $\alpha_{\max}$.)
We need to find a sample cloud of configurations where the Husimi distribution values are substantial, and compare the resulting geometries of Husimi distributions for ladder and non-ladder states. Due to the high dimensionality of the phase space, judicious choices should be made in designing this sampling. If the cloud is not sufficiently spread out, it will overlook the key features of the Husimi distribution in areas falling outside its coverage. (One needs to move away from the maxima of Husimi distributions at $\alpha_{\max}$ to appreciate the extent of their spreading.) If the cloud is too dispersed, an unrealistically large number of sample points will be required in a high-dimensional space to ever hit appreciable values of the Husimi distribution. Thus, we must generate a cloud that restricts the search to a useful region of phase space, without being too restrictive. In practice, after constructing the phase-space point $\alpha_{\max}$ where the Husimi distribution peaks by using optimization algorithms, we use it as the starting value for a Metropolis-Hastings sampling run that generates a cloud of points. To make the procedure effective, as the target probability distribution for this run we choose a Gaussian decaying away from the classical hypersurface (24), reminiscent of the formulas seen in [41, 40],
@@
The subscript ‘$\max$’ refers to the value of the corresponding functional on $\alpha_{\max}$. For concreteness, we focus on a large block of dimension $5604$, with $N=70$ and $M=30$, and restrict our attention to eigenstates with the center-of-mass at rest, annihilated by $Z^{\dagger}$. The relevant ladder state (13) is the lowest energy state in the block. For our choice of a generic state, we pick energy level $5$ above the ground state. To compare the two states, we plot the Husimi function values of the sampled points against their distance $d$ to the two-dimensional surface $\Theta(\alpha_{\max})$. The distance of a phase-space point $\alpha$ to the surface $\Theta(\alpha_{\max})$ is defined as $d=\min_{x\in\Theta(\alpha_{\max})}|x-\alpha|^{2}$. Numerically, we find this distance by drawing a fine grid on $\Theta(\alpha_{\max})$, computing the distance of each point on the grid to $\alpha$. The distance $d$ is the minimum of this
@@
Figure 1: Husimi function values sampled over a cloud of phase-space points, plotted against their distance $d$ to the two-dimensional surface $\Theta(\alpha_{\max})$ where the largest Husimi function value lies. Top: for the ladder state with $|Z|=0$; bottom: for a non-ladder state with $|Z|=0$, both in the Hilbert space block with $N=70$ and $M=30$. The Gaussian distribution used in the sampling algorithm has mean values determined by the classical conservation laws at $\alpha_{\max}$, with $(N_{\max},M_{\max},Z_{\max},H_{\max})=(70,30,0,1925)$ (top) and $(N_{\max},M_{\max},Z_{\max},H_{\max})=(70,30,0,1947.04)$ (bottom), obtained by global optimization of the Husimi distribution starting from a randomly picked $\alpha_{n}$. The standard deviations of the Gaussian distribution used for sampling are set to $5\sqrt{2}\times(\sigma_{N},\sigma_{M},\sigma_{|Z|^{2}},\sigma_{H})=(1,1,N,N)$ for both cases. For the non-ladder state, we further perform local optimization of the Husimi distribution for randomly selected points in the upper right area of the plot, and display the distance from the resulting local maximum to the initial two-dimensional surface, indicated by the color of the selected points. These results show that the Husimi distribution of the ladder state is localized on $\Theta(\alpha_{\max})$, while the region of high Husimi distribution spreads beyond $\Theta(\alpha_{\max})$ for generic states, with many distinct local maxima decorating the landscape.
set. The results are displayed in Fig. 1 for the two chosen states. Some further numerical examples are given in Appendix C, and they are consistent with the general picture presented here. Different choices for the standard deviations in (25) have been tested to obtain an optimal sampling cloud. Clouds that are too dispersed would result in a large sampling of phase-space points with very small $Q_{\psi}$, while clouds that are too restrictive would not sample Husimi function values at distances far enough from $\Theta(\alpha_{\max})$, and would not be able to distinguish localized and non-localized states.
For the ladder state, Fig. 1 shows an exponential correlation between the Husimi value and the distance away from $\Theta(\alpha_{\max})$. This reflects exponential decay away from the classically permitted region of motion, which coincides with $\Theta(\alpha_{\max})$ in this case. On the other hand, for the generic state, the bottom plot in Fig. 1 shows a protrusion of points in the top-right direction where the Husimi function is large while being far from the two-dimensional surface $\Theta(\alpha_{\max})$. By performing a local optimization of $Q_{\psi}$ on phase-space points in this part of the figure, we discover new maxima of $Q_{\psi}$ that are located at significant distances away from $\Theta(\alpha_{\max})$.
VI Conclusions